OpenStudy (anonymous):

On the first 3 exams of his geometry class, Luis got an average score of 84. What does he need on the next exam to have an overall average of 86?

OpenStudy (tkhunny):

\(Average = \dfrac{Total\;of\;ALL\;FOUR\;Scores}{4}\) - You are missing only one piece. Give it a name and dive in!

OpenStudy (anonymous):

it doesn't tell me the scores. Just the average

OpenStudy (tkhunny):

Not directly, but it does tell you. For the first 3 exams, \(Average = \dfrac{Total\;of\;3\;Exams}{3}\implies 84 = \dfrac{What?}{3}\)

OpenStudy (anonymous):

Ah. Ok thanks

OpenStudy (mathstudent55):

The average of the first three scores is 84. That was obtained by dividing the total score of the three tests, and dividing by 3. If you multiply 84 by 3, you'll find the total score of the three tests. Then that total plus the unknown fourth score, all divided by 4 equals 86. Set up an equation and solve for the unknown.

OpenStudy (anonymous):

thank you

OpenStudy (mathstudent55):

You're welcome.